OpenStudy (anonymous):

How do I solve this? x^2/3 + 4x^1/3 + 3 = 0

OpenStudy (anonymous):

Factorize it as : (x^(1/3)+1) (x^(1/3)+3) = 0

OpenStudy (anonymous):

sorry, wrong didn't see the +3

OpenStudy (anonymous):

x never equals zero though... i don't think

OpenStudy (anonymous):

There are no solutions,I suppose

OpenStudy (anonymous):

are you sure you wrote the problem correctly?

OpenStudy (anonymous):

Thanks, I'm still confused...I will retype the question...

OpenStudy (xishem):

How do you figure that x never equals 0?

OpenStudy (anonymous):

graphing calculator

OpenStudy (anonymous):

@ xishem try substituting x = 0

OpenStudy (anonymous):

Solve the equation. x^2/3 + 4x^1/3 + 3 = 0 I think when solving for x, the equation is set equal to zero...

OpenStudy (xishem):

\[f(-1)=0\]

OpenStudy (anonymous):

no, it equals 2

OpenStudy (anonymous):

@studioso33 is it x^(2/3)

OpenStudy (anonymous):

IT has no solutions; plot it and see

OpenStudy (anonymous):

it has no solution as written in the real numbers, there are non-real answers however

OpenStudy (xishem):

\[f(-1)=x^{2/3}+4x^{1/3}+3=(-1)^{2/3}+4(-1)^{1/3}+3=1-4+3=0\]

OpenStudy (anonymous):

SalvatoreTRM, I'm not sure? I've never typed out equations before!

OpenStudy (anonymous):

does it look like Xishem wrote it?

OpenStudy (anonymous):

x= -27 or -1

OpenStudy (xishem):

I plotted, and there are two solutions.

OpenStudy (anonymous):

Sorry, I'm new here...you guys are far more advanced at math than I! Yes, looks like Xishem wrote it...

OpenStudy (anonymous):

ok then yes, Ireneoseaspe is right and Xishem as well. x=-27 and -1

OpenStudy (anonymous):

Xishem, how to I plot and find solutions myself? Are you all using graphing calculators?

OpenStudy (xishem):

Why is Wolfram plotting it like this? http://www.wolframalpha.com/input/?i=x%5E%282%2F3%29%2B4x%5E%281%2F3%29%2B3

OpenStudy (xishem):

@studio: I didn't find the original solutions graphically, but rather by factoring and setting each of the factors equal to zero separately.

OpenStudy (anonymous):

Ok, i take it back, no solutions. Viva wolfram

OpenStudy (anonymous):

Thank you all...it will take time for me to get used to this online tutoring...I will be in here ALOT!

OpenStudy (xishem):

Wolfram is wrong. Take a look at THIS plot, which is just a factorization of the original. http://www.wolframalpha.com/input/?i=%281%2Bx%5E%281%2F3%29%29+%283%2Bx%5E%281%2F3%29%29&lk=1&a=ClashPrefs_*Math- Both the math and my graphing calculator tell me that there are two solutions.

OpenStudy (anonymous):

who is wolfram?

OpenStudy (xishem):

\[(x^{1/3}+1) (x^{1/3}+3) = 0\]\[x^{1/3}+1=0 \rightarrow x^{1/3}=-1 \rightarrow x=(-1)^3 \rightarrow x=-1\]You can't deny that there are solutions.

OpenStudy (anonymous):

@studio website

OpenStudy (anonymous):

If you put x = -1 you get 3 not 0

OpenStudy (anonymous):

x^2/3 + 4x^1/3 + 3 = 0 u=x^1/3 u^2 + 4u + 3 = 0 (u + 3)(u + 1) = 0 u= -3,u= - 1 u=x^1/3 -3 = x^1/3 (-3)^3 = (x^1/3)^3 -27 = x u=x^1/3 -1 = x^1/3 (-1)^3 = (x^1/3)^3 -1 = x

OpenStudy (xishem):

Aron: Show your work. I get 0 every time.

OpenStudy (anonymous):

@Irene if you substitute the values you wont get 0.

OpenStudy (anonymous):

Xishem, do you know robby mayasich?

OpenStudy (anonymous):

nvm, your profile pic reminded me of someone

OpenStudy (xishem):

http://www.wolframalpha.com/input/?i=%28-1%29%5E%281%2F3%29 Wolfram also thinks that is true. Weird.

OpenStudy (xishem):

And that error would explain why WolframAlpha doesn't give the correct answers.

OpenStudy (anonymous):

(x^(1/3)+1) (x^(1/3)+3) = 0

OpenStudy (anonymous):

x^2/3 + 4x^1/3 + 3 = 0 (-27)^2/3 + 4(-27)^1/3 + 3 = 0 \[\sqrt[3]{-27^{2}} + 4\sqrt[3]{-27} + 3 = 0\] \[\sqrt[3]{729} + 4(-3) + 3 = 0\] \[9 -12 + 3 = 0\] \[-3 + 3 =0\] 0=0